# Gravitational Redshift of Light and the Heisenberg Uncertainty Principle

### 30 Mar 2026 v1

@@ Abstract
Keywords: Gravitational redshift of light, Heisenberg uncertainty, photonic entanglement and EPR
@@ I Introduction
Gravitational redshift (GRS) can be derived from first principles and was predicted by Einstein, on the basis of the equivalence principle, prior to the formulation of general relativity (GR) [1]. Subsequent analyses have argued that, unless the energy of a photon traversing a static gravitational potential decreases, the first law of thermodynamics would be violated [^1] The argument proceeds as follows. Consider an excited atom at point 2 with energy, $E_{2}=m_{A}c^{2}+\epsilon=m_{A}^{*}c^{2}$, fall in a gravitational field to a point 1. Gravity does work, $m_{A}^{*}\Delta\Phi$. Let the atom radiate a photon at point 1, $\epsilon=h\nu_{1}$, thereby reducing the atom’s energy to $m_{A}c^{2}$. Now the atom is taken back to point 2. Gravity does work, $m_{A}\Delta\Phi$. If the emitted photon’s frequency is not changed it could excite an identical atom causing a cyclic process in which work is done and energy is produced, $\oint dW=\left(m_{A}^{*}-m_{A}\right)\Delta\Phi=\epsilon\frac{\Delta\Phi}{c^{2}}$. Unless the energy of the photon is decreased by GRS such that $h(\nu_{1}-\nu_{2})=\epsilon\frac{\Delta\Phi}{c^{2}}$, the first law of thermodynamics would be violated [2]. [2]. As a quantum object, any change in the photon’s energy and momentum must comply with the Heisenberg uncertainty principle (HUP). However, cross-verification of the momentum–position uncertainty relation against empirical data reveals a persistent tension. Reformulating this relation solely in terms of wavelength leads to paradoxical constraints that conflict with GR. Furthermore, because photon momentum is not conserved, we develop a model from first-principles that independently substantiates these inconsistencies. Taken together, these results indicate an incompatibility between GR and QM that manifests itself even within the weak-field regime.
@@ II.1 GRS and the Momentum–Position uncertainty
To leading order this reduces to Equation (3b)222:Consider the following, we can recast the position in terms of the wavelength as, $z=q\lambda$, where $q$ is the number of wavelengths that fit into the distance $z$. For one wavelength we have, $\delta z_{1}\sim\delta\lambda\equiv(\delta\lambda/\lambda)\lambda$, with $\delta\lambda/\lambda$, the resolution of the wavelength, usually related to the FWHM. The RMS of these $q$ deviations is, $\Delta z_{\mathrm{rms}}=\sqrt{\sum\frac{1}{q}\left(\frac{\delta\lambda_{i}}{\lambda}\right)^{2}\lambda^{2}}=\left(\frac{\delta\lambda}{\lambda}\right)\lambda$, with a maximum uncertainty, $\Delta z_{\max}=\sqrt{q\left(\frac{\delta\lambda}{\lambda}\right)^{2}\lambda^{2}}=\sqrt{q}\left(\frac{\delta\lambda}{\lambda}\right)\lambda$. The resolution is a numerical factor and the numerical value is renamed as, $\sqrt{q}\left(\frac{\delta\lambda}{\lambda}\right)\equiv\kappa$, leading to $\Delta z=\kappa\lambda$.
@@
Note that the Planck constant, $h$, drops out of Equations (4), underscoring their intrinsically classical character. The wavelength shift of a photon ascending from Earth’s surface to an height $z$ is calculated in GR with the Schwarzschild metric [5]:
@@
$\kappa\geq(2\pi\delta\lambda/\lambda)^{-1}\approx 2.5\times 10^{11}:,$ (7)
@@
The observed agreement with the GR prediction, $\Delta\nu_{\mathrm{obs}}/\Delta\nu_{\mathrm{GR}}\simeq 1$, only sharpens the issue: according to Equation (1), each additional meter of elevation should degrade this accuracy dramatically. What, then, is the operational meaning of the uncertainty principle for a photon wavepacket propagating in Earth’s gravitational field, and how does this comport with the deterministic structure of general relativity? If the photon’s momentum is not a real, well-defined property across its entire wavepacket during its ascent, how can the cumulative effect implied by Equation (5) arise? In this view, spacetime effectively acts as a continuous probe of the photon’s spatial and temporal properties—an assumption that implicitly reinstates a form of local realism. The fact that the measurement procedure respects the uncertainty principle does not resolve the paradox: gravitational redshift appears to rely on a quasi-classical, locally realistic account of the photon’s dynamical variables, in direct tension with the quantum-mechanical prohibition against their simultaneous definiteness.
@@ II.2 GR Imposed Constraints on $\kappa$
The gravitational redshift of a photon emitted from the surface of a massive body at radius $r_{E}$ and observed at earth at a distance $r_{O}\gg r_{E}$, is calculated by the Schwarzschild metric as:
$\frac{\lambda_{O}}{\lambda_{E}}=\frac{\sqrt{1-\frac{2GM}{r_{O}c^{2}}}}{\sqrt{1-\frac{2GM}{r_{E}c^{2}}}}\underset{r_{0}\to\infty}{\cong}\frac{1}{\sqrt{1-\frac{2GM}{r_{E}c^{2}}}}.$ (9)
The quantity $\epsilon=\frac{2GM}{r_{E}c^{2}}$ is the ratio between the Schwarzschild radius, $\frac{2GM}{c^{2}}$, and the physical radius of the star, $r_{E}$. The fractional shift is therefore:
@@
GRS values increase as $\epsilon$ approaches the size of its black hole radius. Indeed as $\epsilon\rightarrow 1$, the shift becomes infinite. An infinite shift is consistent with the fact that light cannot escape the horizon of a black hole.
In the following, we shall assume that the numerical values predicted by Equation (9), as they may be realized through a quantum measurement protocol, are typically smaller than the intrinsic resolution set by the linewidth $\Gamma$. That is, for a measurement procedure involving large statistics, gravity shifts the center of the distribution, whereas the linewidth reflects the underlying quantum uncertainty. However, for the interaction of a single photon with spacetime, the gravitationally induced shift would be obscured by this intrinsic uncertainty. Under the assumption that gravity acts on individual photons according to GR, we express the relation between gravitational redshift and the Heisenberg uncertainty principle as:
@@
$\kappa(\epsilon)\geq\frac{\sqrt{1-\epsilon}}{4\pi(1-\sqrt{1-\epsilon})}.$ (12)
@@
Figure 1: GRS as calculated from Equation 9. The insert is a log-log plot of Equation 11.
@@ III HUP and Entanglement
Photonic entanglement deems quantum mechanics a nonlocal theory, in contradiction with general relativity and the equivalence principle. This is best described by the EPR paradox [6]. Entangled photons that undergo gravitational redshift would entangle spacetime itself, thereby turning it into a quantum entity. So far, as has been observed in numerous tests of general relativity, spacetime adheres to its local character. Extending the argument of local realism from single quanta photons to entangled photons would mean that the entangled pair violates the uncertainty principle with respect to the independent reality of their positions and momenta in spacetime. No “steering” [^3] Schrödinger introduced the concept of entanglement or “steering” to express the nonlocal effect of part of a system on another part of it, in what seems to be an action at a distance. information has to travel through spacetime faster than the speed of light in order to maintain the correlation between the members of the pair. The initial correlation would be altered by the local interaction of each photon of the pair with spacetime. As no measurement is performed the “collapse of the wavefunction” is avoided altogether.
@@
To this end we propose to test the effect with a thought experiment utilizing a set of EPR entangled beams on Earth’s surface. Entanglement-based EPR arguments presuppose a single global time parameter $t$ with respect to which joint observables such as:
$\widehat{X}_{-}=\widehat{x}_{A}(t)-\widehat{x}_{B}(t)$
and
$\widehat{P}_{+}=\widehat{p}_{A}(t)+\widehat{p}_{B}(t)$
are defined as simultaneous measurements on spatially separated subsystems. In a curved spacetime, by contrast, each subsystem follows a worldline, $x_{j}^{\mu}(\tau_{j})$, with its own proper time, $\tau_{j}$, related to any chosen coordinate time $t$ by:
@@
which in the weak-field limit reduces to:
$d\tau_{j}\simeq\left(1+\Phi_{j}/c^{2}\right)dt.$
The local mode operators evolve as:
@@
$(\widehat{x}_{j},\widehat{p}_{j})\mapsto(\sqrt{\zeta_{j}}\,\widehat{x}_{j},\,\widehat{p}_{j}/\sqrt{\zeta_{j}})$
with $\zeta_{j}\simeq\left(1+\Phi_{j}/c^{2}\right)$.
The global EPR observables $\widehat{X}_{-}$ and $\widehat{P}_{+}$ become explicitly gravity-dependent constructs whose canonical form is not invariant under changes of gravitational potential, even though the underlying bipartite state should remain entangled [^4] This issue has, in fact, already manifested in the discussion of a single Mössbauer wavepacket (Sec. 2-A).. Thus, the mathematical structure of general relativity—through the replacement of a universal time $t$ by different proper times $\tau_{A},\tau_{B}$—undermines the very notion of globally defined, simultaneously measurable EPR observables, exposing a tension between the EPR state and dilation of time in a weak gravitational field. Moreover, the entanglement of the $\lvert\mathrm{EPR}\rangle$ state eliminates the uncertainty of the combined variables, $\widehat{X}_{-}$ and $\widehat{P}_{+}$. Specifically:
@@ III.1 Proposed Experiment
As gravitational redshift offers the stage on which quantum mechanics is pitted against general relativity we propose to use photonic entanglement as an experimental testbed of local realism per EPR. In quantum mechanics a pair of entangled photons is viewed as a single system which maintains its correlation as long as it is not subject to collapse due to measurement. Even if the system undergoes decoherence due to gravity the lifetime of the correlation between the photons of the pair can be estimated by the coherence length:
@@
Should the entangled pair obey the uncertainty relation, then, within the time estimated above, the correlation between the members of the pair should stay unperturbed. Hence, by testing if entanglement withstand gravitational redshift, or within what time it breaks down, we should be able to test the effect of a gravitational field on a quantum system. As depicted schematically in figure 2, Alice would be observing at ground level the interference pattern that arises between the laser beams A and A’, both with the same wavelengths.
@@
While the first three outcomes indicate that entanglement influences spacetime and may therefore require incorporation into general relativity, the fourth outcome instead suggests that quantum action is fundamentally local and that GR does not require quantum degrees of freedom at all.
@@
Finally, it should also be noted that the laser beams can be polarized, which makes it possible to perform Bell-inequality measurements concurrently with the continuous-variable entanglement at both Alice’s and Bob’s stations. The results of these measurements would make it possible to compare the nature of these two completely different forms of entanglement.
@@
By polarizing the beams, it becomes possible to perform Bell-inequality measurements concurrently with the continuous-variable entanglement at both ends of the setup. The results of these measurements would make it possible to compare the nature of these two fundamentally different forms of entanglement.
@@ Acknowledgements
Asher Klatchko gratefully acknowledges David Griffiths, Emeritus Professor of Physics at Reed College, Robert Hamilton of George Fox University, Sandu Popescu of Bristol University and Daniel Rubén Zerzión of DIPC San Sebastián for insightful discussions and valuable comments.

### 01 May 2026 v2

@@ Abstract
Keywords: Gravitational redshift of light, Heisenberg uncertainty, photonic entanglement vis-à-vis EPR
@@ I Introduction
Gravitational redshift (GRS) can be derived from first principles and was predicted by Einstein, on the basis of the equivalence principle, prior to the formulation of general relativity (GR) [1]. Subsequent analyses have argued that, unless the energy of a photon traversing a static gravitational potential decreases, the first law of thermodynamics would be violated [^1] The argument proceeds as follows. Consider an excited atom at point 2 with energy, $E_{2}=m_{A}c^{2}+\epsilon=m_{A}^{*}c^{2}$, falling in a gravitational field to a point 1. Gravity does work, $m_{A}^{*}\Delta\Phi$. Let the atom radiate a photon at point 1, $\epsilon=h\nu_{1}$, thereby reducing the atom’s energy to $m_{A}c^{2}$. Now the atom is taken back to point 2. Gravity does work, $m_{A}\Delta\Phi$. If the emitted photon’s frequency is not changed it could excite an identical atom causing a cyclic process in which work is done and energy is produced, $\oint dW=\left(m_{A}^{*}-m_{A}\right)\Delta\Phi=\epsilon\frac{\Delta\Phi}{c^{2}}$. Unless the energy of the photon is decreased by GRS such that $h(\nu_{1}-\nu_{2})=\epsilon\frac{\Delta\Phi}{c^{2}}$, the first law of thermodynamics would be violated [2]. [2]. Note however that in the covariant formalism of GR, the actual conservation law in any static spacetime with an independent metric satisfying $g_{0j}=0$ is,
$E=-p_{0}=-g_{00}p^{0}=|g_{00}|^{1/2}\,E_{\text{local}}=\text{constant}.$
This expresses the law of energy redshift, describing how the locally measured energy of any particle—massless or massive—changes as it "climbs out" of or "falls into" a static gravitational field. With $E_{\text{local}}=h~c/\lambda_{\text{local}}$, this conservation law becomes [^2] See for example Equation (8) below. :
$|g_{00}|^{-1/2}\,\lambda_{\text{local}}=\text{constant}.$
Thus, it is not the local energy itself that is conserved between different observers, but rather the above combination — the energy at infinity in terms of these locally measured quantities [3]. Hence, for a photon moving freely from a given observer to infinity, the correction factor reduces to unity and the local energy (as measured at infinity) becomes identical with the constant E.
As a quantum object, any change in the photon’s energy and momentum must comply with the Heisenberg uncertainty principle (HUP). However, cross-verification of the momentum–position uncertainty relation against empirical data reveals a persistent tension. Reformulating this relation solely in terms of wavelength leads to paradoxical constraints that conflict with GR. Furthermore, because photon momentum is not conserved, we develop a model from first-principles that independently substantiates these inconsistencies. Taken together, these results indicate an incompatibility between GR and QM that manifests itself even within the weak-field regime.
@@ II.1 GRS and the Momentum–Position uncertainty
To leading order this reduces to Equation (3b) [^3] Consider the following, we can recast the position in terms of the wavelength as, $z=q\lambda$, where $q$ is the number of wavelengths that fit into the distance $z$. For one wavelength we have, $\delta z_{1}\sim\delta\lambda\equiv(\delta\lambda/\lambda)\lambda$, with $\delta\lambda/\lambda$, the resolution of the wavelength, usually related to the FWHM. The RMS of these $q$ deviations is, $\Delta z_{\mathrm{rms}}=\sqrt{\sum\frac{1}{q}\left(\frac{\delta\lambda_{i}}{\lambda}\right)^{2}\lambda^{2}}=\left(\frac{\delta\lambda}{\lambda}\right)\lambda$, with a maximum uncertainty, $\Delta z_{\max}=\sqrt{q\left(\frac{\delta\lambda}{\lambda}\right)^{2}\lambda^{2}}=\sqrt{q}\left(\frac{\delta\lambda}{\lambda}\right)\lambda$. The resolution is a numerical factor and the numerical value is renamed as, $\sqrt{q}\left(\frac{\delta\lambda}{\lambda}\right)\equiv\kappa$, leading to $\Delta z=\kappa\lambda$.
@@
Note that the Planck constant, $h$, drops out of Equations (4), underscoring their intrinsically classical character.
The wavelength shift of a photon ascending from Earth’s surface to an height $z$ is calculated in GR with the Schwarzschild metric [5]:
@@
$\kappa\geq(2\pi\delta\lambda/\lambda)^{-1}\approx 2.5\times 10^{11},$ (7)
@@
The observed agreement with the GR prediction, $\Delta\nu_{\mathrm{obs}}/\Delta\nu_{\mathrm{GR}}\simeq 1$, only sharpens the issue: according to Equation (1), each additional meter of elevation should dramatically degrade this accuracy. What, then, is the operational meaning of the uncertainty principle for a photon wavepacket propagating in Earth’s gravitational field, and how does this comport with the deterministic structure of general relativity? If the photon’s momentum is not a real, well-defined property across its entire wavepacket during its ascent, how can the cumulative effect implied by Equation (5) arise? In this view, spacetime effectively acts as a continuous probe of the photon’s spatial and temporal properties—an assumption that implicitly reinstates a form of local realism. The fact that the measurement procedure respects the uncertainty principle does not resolve the paradox: gravitational redshift appears to rely on a quasi-classical, locally realistic account of the photon’s dynamical variables, in direct tension with the quantum-mechanical prohibition against their simultaneous definiteness.
@@ II.2 GR Imposed Constraints on $\kappa$
The gravitational redshift of a photon emitted from the surface of a massive body at radius $r_{E}$ and observed at earth at a distance $r_{O}\gg r_{E}$, is calculated by the Schwarzschild metric as follows:
$\frac{\lambda_{O}}{\lambda_{E}}=\frac{\sqrt{1-\frac{2GM}{r_{O}c^{2}}}}{\sqrt{1-\frac{2GM}{r_{E}c^{2}}}}~~r_{0}\to\infty~~\frac{1}{\sqrt{1-\frac{2GM}{r_{E}c^{2}}}}.$ (9)
The quantity $\epsilon=\frac{2GM}{r_{E}c^{2}}$ is the ratio between the Schwarzschild radius, $\frac{2GM}{c^{2}}$, and the physical radius of the star, $r_{E}$. The fractional shift is, therefore:
@@
GRS values increase as $\epsilon$ approaches the size of its black hole radius. Indeed, as $\epsilon\rightarrow 1$, the shift becomes infinite. An infinite shift is consistent with the fact that light cannot escape the horizon of a black hole.
In the following, we shall assume that the numerical values predicted by Equation (9), as may be realized through a quantum measurement protocol, are typically smaller than the intrinsic resolution set by the linewidth $\Gamma$. That is, for a measurement procedure involving large statistics, gravity shifts the center of the distribution, whereas the linewidth reflects the underlying quantum uncertainty. However, for the interaction of a single photon with spacetime, the gravitationally induced shift would be obscured by this intrinsic uncertainty. Under the assumption that gravity acts on individual photons according to GR, we express the relation between gravitational redshift and the Heisenberg uncertainty principle as:
@@
$\kappa(\epsilon)\geq\frac{\sqrt{1-\epsilon}\,\bigl(1+\sqrt{1-\epsilon}\bigr)}{4\pi\epsilon}\sim\frac{1}{2\pi\epsilon}.$ (12)
@@
[…]
Figure 1: (a) GRS as calculated from Equation (9). (b) a log-log plot of Equation (11).
@@ III HUP and Entanglement
Photonic entanglement deems quantum mechanics a nonlocal theory, in contradiction with general relativity and the equivalence principle. This is best described by the EPR paradox [7]. Entangled photons that undergo gravitational redshift would entangle spacetime itself, thereby turning it into a quantum entity. So far, as has been observed in numerous tests of general relativity, spacetime adheres to its local character. Extending the argument of local realism from single quanta photons to entangled photons would mean that the entangled pair violates the uncertainty principle with respect to the independent reality of their positions and momenta in spacetime. No “steering” [^4] Schrödinger introduced the concept of entanglement or “steering” to express the nonlocal effect of part of a system on another part of it, in what seems to be an instantaneous action at a distance. information has to travel through spacetime faster than the speed of light in order to maintain the correlation between the members of the pair. The initial correlation would be altered by the local interaction of each photon of the pair with spacetime. As no measurement is performed the “collapse of the wavefunction” is avoided altogether.
@@
To this end we propose to test the effect with a thought experiment utilizing a set of EPR entangled beams on Earth’s surface. Entanglement-based EPR arguments presuppose a single global time parameter $t$ with respect to joint observables such as:
$\widehat{X}_{-}=\widehat{x}_{A}(t)-\widehat{x}_{B}(t)~\text{and}~\widehat{P}_{+}=\widehat{p}_{A}(t)+\widehat{p}_{B}(t),$
which are defined as simultaneous measurements on spatially separated subsystems. In a curved spacetime, by contrast, each subsystem follows a worldline, $x_{j}^{\mu}(\tau_{j})$, with its own proper time, $\tau_{j}$, related to any chosen coordinate time $t$ by:
@@
which in the weak-field limit reduces to, $d\tau_{j}\simeq\left(1+\Phi_{j}/c^{2}\right)dt$. The local mode operators evolve as:
@@
$(\widehat{x}_{j},\widehat{p}_{j})\mapsto(\sqrt{\zeta_{j}}\,\widehat{x}_{j},\,\widehat{p}_{j}/\sqrt{\zeta_{j}})~\text{with}~\zeta_{j}\simeq\left(1+\Phi_{j}/c^{2}\right).$
The global EPR observables $\widehat{X}_{-}$ and $\widehat{P}_{+}$ become explicitly gravity-dependent constructs whose canonical form is not invariant under changes of gravitational potential, even though the underlying bipartite state should remain entangled [^5] This issue has, in fact, already manifested in the discussion of a single Mössbauer wavepacket (Sec. II.A).. Thus, the mathematical structure of general relativity—through the replacement of a universal time $t$ by different proper times $\tau_{A},\tau_{B}$—undermines the very notion of globally defined, simultaneously measurable EPR observables, exposing a tension between the EPR state and dilation of time in a weak gravitational field. Moreover, the entanglement of the $\lvert\mathrm{EPR}\rangle$ state eliminates the uncertainty of the combined variables, $\widehat{X}_{-}$ and $\widehat{P}_{+}$. Specifically:
@@ III.1 Proposed Experiment
As gravitational redshift offers the stage on which quantum mechanics is pitted against general relativity we propose to use photonic entanglement as an experimental testbed of local realism per EPR. In quantum mechanics a pair of entangled photons is viewed as a single system which maintains its correlation as long as it is not subject to collapse due to measurement. Even if the system undergoes (quantum) decoherence [11] due to gravity  [^6] Induced gravitational quantum phase shifts were first observed in neutron‑interference experiments by Colella, Overhauser, and Werner [12]. Likewise, massless photons in curved spacetime acquire a phase, $\Phi=\frac{1}{\hbar}\int p_{\mu}\,dx^{\mu}$, determined by the underlying geometry [13]. Decoherence does not alter the phase itself; rather, it governs whether quantum states can interfere or become distinguishable. Within this viewpoint, the system could be regarded as becoming entangled with quantum degrees of freedom attributed to spacetime, leading to the suppression of interference and the loss of off‑diagonal density‑matrix elements when those degrees of freedom are traced out. However, any such attribution remains speculative: the existence and nature of these putative degrees of freedom would depend on a theory of quantum gravity that is not currently established. Consequently, this interpretation should be regarded with appropriate caution. the lifetime of the correlation between the photons of the pair can be estimated by the coherence length:
@@
Should the entangled pair obey the uncertainty relation, then, within the time estimated above, the correlation between the members of the pair should stay unperturbed. Hence, by testing if entanglement withstands gravitational redshift, or within what time it breaks down, we should be able to test the effect of a gravitational field on a quantum system. More precisely, we seek to determine whether the entanglement of the pair enforces the conservation of the system’s initial energy after one of its members has traveled a height $z$:
$\Delta E_{\text{pair}}=\hbar(\omega^{(A)}_{h=0}-\omega^{(B)}_{h=0})\;\overset{?}{=}\;\hbar(\omega^{\text{(A)}}_{h=0}-\omega^{\text{(B)}}_{h=z})$.
If the initial energy of the entangled EPR state were shown to remain conserved after one constituent has been displaced to a height $z$, such conservation would necessarily be attributable solely to the underlying entanglement structure. This result would, in turn, provide evidence that quantum‑mechanical degrees of freedom play an essential role in gravitational contexts and therefore must be incorporated into the theoretical framework of general relativity.
As depicted schematically in figure 2, Alice would be observing at ground level the interference pattern that arises between the laser beams A and A’, both with the same wavelengths.
@@
While the first three outcomes indicate that entanglement influences spacetime and may therefore require incorporation into general relativity, the fourth outcome instead suggests that the photon interaction with spacetime is fundamentally local and that GR does not require quantum degrees of freedom at all.
@@
Because the beams are polarization entangled, Bell‑inequality tests can be performed concurrently with the continuous‑variable entanglement. A polarized, entangled‑beam arrangement analogous to Ref. [12] would offer the same capability. Appendix A sketches the underlying theory.
@@ Acknowledgements
Asher Klatchko gratefully acknowledges David Griffiths, Emeritus Professor of Physics at Reed College, Robert Hamilton of George Fox University, Sandu Popescu of Bristol University and Daniel Rubén Zerzión of DIPC San Sebastián for insightful discussions and valuable comments. We also wish to acknowledge productive comments by Germano D’Abramo from Ministero dell’Istruzione, Università e Ricerca.
Appendix A Bell type experiment with a gravitational induced phase shift
This appendix contrasts gravitationally induced phase shifts in polarization entanglement with the field‑dependent energy effects in continuous‑variable entanglement.
Consider a polarization-entangled photon pair prepared in the Bell state
$|\Psi_{0}\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle_{A}|H\rangle_{B}+|V\rangle_{A}|V\rangle_{B}\right).$ (18)
Photon $A$ propagates along a horizontal reference path, while photon $B$ follows a trajectory $\gamma$ in a weak, static gravitational field. In the weak-field limit, the metric is
$ds^{2}=-\left(1+\frac{2gz}{c^{2}}\right)c^{2}dt^{2}+d\mathbf{x}^{2}.$ (19)
The phase accumulated by a photon along $\gamma$ is
$\Phi=\int_{\gamma}k_{\mu}dx^{\mu},$ (20)
where $k^{\mu}$ is the null wavevector. In a stationary spacetime, the conserved quantity
$\omega=-k_{\mu}\xi^{\mu}$ (21)
(with $\xi^{\mu}=(\partial_{t})^{\mu}$) defines the photon frequency. To leading order in $gz/c^{2}$, the phase difference between two paths reduces to
$\Phi=\omega\Delta\tau,$ (22)
where the effective proper-time difference is
$\Delta\tau=\frac{1}{c^{2}}\int gz(t)\,dt.$ (23)
Thus,
$\Phi=\frac{\omega}{c^{2}}\int gz(t)\,dt.$ (24)
After propagation, the entangled state becomes
$|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|H\rangle_{A}|H\rangle_{B}+e^{i\Phi}|V\rangle_{A}|V\rangle_{B}\right).$ (25)
This phase is observable in joint polarization measurements. In the basis
$|\pm\rangle=\frac{1}{\sqrt{2}}(|H\rangle\pm|V\rangle),$ (26)
the coincidence probability reads
$P_{++}=P_{--}=\frac{1}{2}\left(1+\cos\Phi\right)~\text{and}$ (27)
$P_{+-}=P_{-+}=\frac{1}{2}\left(1-\cos\Phi\right).$ (28)
As Eq.(A7) shows, the gravitational field induces a Bell‑detectable relative phase that, depending on the measurement basis, can at most obscure—but not eliminate—the entanglement, whereas the redshift in Eq.(1) can destroy the underlying momentum‑entanglement structure altogether.
[…]
Figure 3: Schematic of an entangled–photon interferometer in which Photon A follows a horizontal reference path while Photon B traverses a vertical interferometer and acquires a gravitationally induced phase $\Phi$. The phase shift is read out through polarization–correlation measurements in coincidence.